# The Essence of Contextual Understanding in Theory of Mind: A Study on Question Answering with Story Characters

### v1

@@ Abstract
Theory-of-Mind (ToM) is a fundamental psychological capability that allows humans to understand and interpret the mental states of others. Humans infer others’ thoughts by integrating causal cues and indirect clues from broad contextual information, often derived from past interactions. In other words, human ToM heavily relies on the understanding about the backgrounds and life stories of others. Unfortunately, this aspect is largely overlooked in existing benchmarks for evaluating machines’ ToM capabilities, due to their usage of short narratives without global backgrounds. In this paper, we verify the importance of understanding long personal backgrounds in ToM and assess the performance of LLMs in such realistic evaluation scenarios. To achieve this, we introduce a novel benchmark, CharToM-QA, comprising 1,035 ToM questions based on characters from classic novels. Our human study reveals a significant disparity in performance: the same group of educated participants performs dramatically better when they have read the novels compared to when they have not. In parallel, our experiments on state-of-the-art LLMs, including the very recent o1 model, show that LLMs still perform notably worse than humans, despite that they have seen these stories during pre-training. This highlights the limitations of current LLMs in capturing the nuanced contextual information required for ToM reasoning. [^1] Data for evaluation will be released after review.
@@ 1 Introduction
Recently, a range of emerging Large Language Models (LLMs) have shown remarkable performance in solving complex tasks and generating human-like language. To further enhance LLMs’ ability to understand user requirements, it is fascinating to explore the extent to which LLMs capture the ToM capabilities and ultimately to equip them with human-level ToM. To this end, researchers have developed numerous benchmarks to evaluate the ToM capabilities of LLMs Grant et al. (2017); Nematzadeh et al. (2018); Sap et al. (2019); Le et al. (2019); Gandhi et al. (2023); Kosinski (2023); Zhou et al. (2023a); Wu et al. (2023); Shapira et al. (2024); Chen et al. (2024). These benchmarks simulate the experimental design in psychology or cognitive research Wimmer and Perner (1983); Dyck et al. (2001), and contain elaborately crafted, story-based tasks that target ToM dimensions originally defined in psychological research.
@@
Despite their comprehensive coverage of ToM dimensions, these existing benchmarks all largely overlook the importance of comprehensive contextual understanding in human ToM. In daily lives, humans understand the minds of others based on the long historical contexts of their complex social relationships and interactions, rather than with only the local circumstances on the spot. This is a natural phenomenon as the behaviors and mental states of a person are greatly influenced by other global background factors, such as their personality and past experience. However, the testing scenarios in existing ToM datasets are usually short of capturing these factors, as they are typically depicted with brief text pieces consisting of only superficial character actions and surrounding environment.
@@
In this paper, we propose CharToM-QA benchmark that evaluates the capability of LLMs on understanding ToM of characters in famous novel books. In CharToM-QA, the task takes the form of ToM-related question answering about characters within story plots. This setting naturally addresses the aforementioned challenges of most existing datasets due to the intrinsic features of story plots in novel books: 1) diverse social scenarios; 2) rich in complex social relationships and interactions; 3) high relevance to the whole book storyline. Thus, it alleviates heavy reliance on pre-determined rules to generate testing scenarios and raises higher requirements for comprehensively understanding context when evaluating ToM capability of current LLMs.
@@
For the dataset creation, we adopt an AI-assisted human annotation strategy that exploits publicly available book notes written by online reading app users. As users read, they can underline a specific text fragment and write down their comments as book notes. These book notes are often closely related to the selected fragments and their surrounding plots, which contain valuable interpretation, analysis and thoughts of human readers. We first collect a large number of user notes and their linked text from novel books. Then, as in most previous ToM assessments Nematzadeh et al. (2018); Sap et al. (2019); Tracey et al. (2022); Gandhi et al. (2023); Ma et al. (2023); Wu et al. (2023); Zhou et al. (2023b), we keep the notes reflecting certain ToM dimensions of appointed characters, including belief, intention, emotion and desire. Based on these notes, GPT-4o is leveraged to generate ToM descriptions of the four dimensions about given characters. Afterwards, these descriptions are used to construct ToM-related questions with answers. During the whole process, strict validation procedures are carried out by experts with strong literature background to ensure data quality.
@@
For evaluation, both generative and multichoice question answering (QA) are conducted. In generative QA, models are asked to give their responses given a question and its corresponding story plot. The qualities of responses are assessed based on the annotated answers. Particularly, traditional token-based metrics (e.g. Rouge Lin (2004)) and embedding-based metrics (e.g. cosine similarity between sentence embeddings) cannot give a fine-grained assessment that reflects how well a response matches the answer. Therefore, we design an evaluation protocol that mimics the process of grading papers. Specifically, we extract critical points from the answers, which are expected to be included in responses as bonus points. Then, GPT-4o is used as evaluator for assessment based on these bonus points. Besides, the evaluator also serves like a criticizer to point out defects existing in responses as penalty if there is any inappropriate statement. The bonus point coverage and penalty rate are used to comprehensively assess the qualities of model responses. In multichoice QA, we elaborately construct distraction choices and measure the model accuracy.
@@
We experiment with several current popular LLMs in both generative QA and multichoice QA settings, where different lengths of plot window are exposed to models. The experimental results show that GPT-4o consistently outperforms other LLMs across the four ToM dimensions in both settings. Moreover, human studies are also conducted using multichoice QA. The results reveal that (1) Humans who have read the book greatly outperform the state-of-the-art LLM, while those who have not perform only on par with it. (2) Human ToM largely benefits from familiarity with the novels. (3) Humans achieve higher accuracy after reading longer contexts. In contrast, the performances of LLMs almost stay stable with different lengths of plot window. This indicates that the ToM capability of current LLMs still struggles at dealing with complex scenarios and capturing nuanced historical contextual information for robust ToM comprehension.
@@ 2 Related Work
With the advent of LLMs, many benchmarks have been developed to evaluate the ToM capability of LLMs by simulating the cognitive experiments originally in psychology Grant et al. (2017); Nematzadeh et al. (2018); Sap et al. (2019); Le et al. (2019); Gandhi et al. (2023); Kosinski (2023); Zhou et al. (2023a); Wu et al. (2023); Shapira et al. (2024); Chen et al. (2024). Le et al. (2019) proposed an improved evaluation protocol and dataset by carefully examining the answer space to control for data regularities to mitigate dataset bias. Sap et al. (2019) introduced Social IQa that tests social and emotional intelligence, requiring models to infer people’s intentions and emotions. Wu et al. (2023) explored higher-order ToM that involves recursive reasoning on others’ beliefs. Tom-in-Mac Yu et al. (2024) assessed meta-learning capability of machine ToM by designing the task of few-shot character understanding. Chen et al. (2024) introduced ToMBench that encompasses 8 tasks and 31 abilities in social cognition and extensively evaluated popular LLMs across tasks and abilities.
Despite existing ToM benchmarks have covered most dimensions defined in psychology, the constructing procedures and testing scenarios of these datasets are still hindered by some inherent limitations. 1) The story generation procedures of these benchmarks rely heavily on pre-determined rules and templates Nematzadeh et al. (2018); Le et al. (2019); Wu et al. (2023), which constrains it from simulating diverse scenarios in reality. This also increases the risk of incorporating predictable regularities and spurious correlations into datasets, bringing about so-called “Clever Hans” phenomenon Lapuschkin et al. (2019). 2) The testing scenarios in these datasets are usually short of complex social relationships and interactions so that the importance of comprehensive contextual understanding is almost ignored.
3 Problem Definition of ToM in Global Contexts
@@
We consider a set of ToM dimensions prevalently studied in previous work in Section 2, Belief, Intention, Emotion and Desire. The importance of these dimensions in daily lives have been discussed in Apperly (2010). We follow the standard definitions of these dimensions in Apperly (2010), with the detailed definitions provided in Appendix B.
@@ 4 CharToM-QA: Assessing the Contextual Aspect of ToM
We create the CharToM-QA benchmark, designed to assess the ToM capability of LLMs in understanding characters in novel books. The benchmark consists of questions about the minds of characters within story plots in four ToM dimensions. The question and answers are annotated by four human experts in literature, with an AI-assisted annotation strategy that leverages massive publicly available book notes from online reading platforms.
The dataset construction undergoes three steps: (1) book notes collection & filtering; (2) key note extraction & paraphrasing; (3) question generation & verification. Our dataset supports evaluation in both generative QA and multichoice QA settings.
@@ 4.1 Book Notes Collection & Filtering
Next, we filter the notes and keep those that can express any one of the four ToM dimensions of characters within story plots. The prompt used for note filtering is detailed in Appendix D. To ensure the validity of these filtered notes, we also manually check their correctness. Specifically, we sample 100 notes from four books with two annotators who are familiar with the books. For each note, the annotators are required to judge whether it appropriately describes the ToM of the appointed character. The results show that 1) 80% notes are judged as accurate; 2) 13% of the notes are open to subjective interpretation by different readers, yet they remain consistent with the book’s content and cannot be falsified; 3) only 7% notes are inappropriate, which demonstrates the overall validity of our filtered notes. Overall, 93% of notes can be treated as accurate.
@@ 4.2 Construction of Generative QA Task
At the extraction stage, the annotators are instructed to extract the critical part of a note that explicitly reflects one of the four ToM dimensions of a target character, which we call as “key note”. The basic principle of this stage is that annotators should only consider the literal meaning of a note with no additional inference. We elaborately craft examples for every ToM dimension and write a detailed manual to guide annotators. An annotator is only considered qualified after passing a labeling trial. They are given necessary information linked with a user note, including the target character, the underlined text and its surrounding context. After extraction, the key notes are kept for each ToM dimension. Nevertheless, the extracted key notes are often in a form of fragmented expressions that cannot be immediately used as complete ToM descriptions. Thus, a paraphrasing stage is required to convert fragmented parts into complete statements. For each note, we prompt GPT-4o to conduct paraphrasing based on key notes. The details of our used prompt template are presented in Appendix E. Annotators also check the paraphrased descriptions and make minor edit if necessary.
@@
In this step, we aim to construct ToM-related questions using those ToM descriptions. Figure 3 depicts the process of question generation. Particularly, GPT-4o serves as a question generator to produce a set of candidate questions asking about target characters. During generation, along with the ToM description, we also provide GPT-4o with corresponding underlined text and its surrounding context of current story plot. For each ToM description, GPT-4o is asked to generate several candidate questions that adhere to the following requirements:
@@
Subsequently, a verification procedure is involved to preliminarily filter out inappropriate candidates that violate the above requirements. We find that it is better not to let GPT-4o be aware of contextual information about the story plot. This is because GPT-4o tends to build trivial connections between question and answer using any possible detail in the story regardless of importance, preserving too many low-quality candidates. The prompt templates used for generation and verification are given in Appendix F. In this way, we keep at most 4 candidate questions for each ToM description. Then, annotators are responsible for choosing the best one out of these candidates, where slight modifications by human are allowed during examination.
@@ 4.3 Construction of Multichoice QA Task
Besides the aformentioned Generative QA setting, we support multichoice QA evaluation with elaborately constructed distraction choices for each question. Given the question-answer pair within a story plot, GPT-4o is used to produce distraction choices that are seemingly plausible but actually unreasonable. Appendix H provides the details of constructing such difficult distraction choices. Finally, a question is paired with four choices including the correct answer and three incorrect ones, as shown in the Appendix K.
@@ 5 Evaluation Protocols for Generative QA
Until now, we have acquired ToM-related question-answer pairs along with their linked story plots, which can be used to conduct evaluation in a generative QA setting. Models are given story plots and asked questions to make responses whose qualities are assessed using corresponding answers as reference. Conventionally, token-based metrics (e.g. Rouge Lin (2004)) and embedding-based metrics (e.g. cosine similarity between sentence embeddings) are mainly used for the assessment of generative QA tasks. However, these two types of metrics have their inherent limitations. For token-based metrics, they cannot deal with the cases where 1) same meanings are expressed with totally different tokens; 2) non-critical tokens dominate in a sentence. Meanwhile, embedding-based metrics only coarsely measure the semantic similarity but cannot explicitly indicate the pros and cons of a response.
Therefore, inspired by the process of grading papers, we design an evaluation protocol that inspects the bonus points and penalty of a model response. Specifically, for each question, we extract the critical points of its reference answer with the assistance of GPT-4o, which are expected to be included as bonus points in a response. The core requirements for bonus point extraction are: 1) Bonus points must be derived from reference answers with no hallucination; 2) Different bonus points should orient to different aspects of an answer. The statistics about the bonus points of questions in each dimension is given in Appendix C. During evaluation, a GPT-4o evaluator measures the coverage of bonus points as an indicator of response quality. Moreover, it also criticizes the defects in a response as penalty if there is any inappropriate statement. In this way, both bonus point coverage and penalty rate are used as a comprehensive assessment of response quality. The prompts for extracting bonus points and conducting such evaluation protocol are detailed in Appendices G and J, respectively. Appendix K also gives illustrative cases of such evaluation.

### v2

@@ Abstract
Theory-of-Mind (ToM) is a fundamental psychological capability that allows humans to understand and interpret the mental states of others. Humans infer others’ thoughts by integrating causal cues and indirect clues from broad contextual information, often derived from past interactions. In other words, human ToM heavily relies on the understanding about the backgrounds and life stories of others. Unfortunately, this aspect is largely overlooked in existing benchmarks for evaluating machines’ ToM capabilities, due to their usage of short narratives without global context, especially personal background of characters. In this paper, we verify the importance of comprehensive contextual understanding about personal backgrounds in ToM and assess the performance of LLMs in such complex scenarios. To achieve this, we introduce CharToM-QA benchmark, comprising 1,035 ToM questions based on characters from classic novels. Our human study reveals a significant disparity in performance: the same group of educated participants performs dramatically better when they have read the novels compared to when they have not. In parallel, our experiments on state-of-the-art LLMs, including the very recent o1 and DeepSeek-R1 models, show that LLMs still perform notably worse than humans, despite that they have seen these stories during pre-training. This highlights the limitations of current LLMs in capturing the nuanced contextual information required for ToM reasoning.  [^1] Our data can be downloaded at Github https://github.com/Encyclomen/CharToM-QA and Huggingface https://huggingface.co/datasets/ZeroXeno/CharToM-QA.
@@ A Study on Question Answering with Story Characters
Figure 1: An example from CharToM-QA benchmark. LLMs make their responses given a question and corresponding story plot in generative and multichoice QA settings.
@@ 1 Introduction
Recently, a range of emerging Large Language Models (LLMs) have shown remarkable performance in solving complex tasks and generating human-like language. To further enhance LLMs’ capability to understand human, it is fascinating to explore the extent to which LLMs capture the ToM capabilities and ultimately to equip them with human-level ToM. To this end, researchers have developed numerous benchmarks to evaluate the ToM capabilities of LLMs Grant et al. (2017); Nematzadeh et al. (2018); Sap et al. (2019); Le et al. (2019); Gandhi et al. (2023); Wu et al. (2023); Kim et al. (2023); Kosinski (2023); Zhou et al. (2023); Shapira et al. (2024); Xu et al. (2024a); Chan et al. (2024); Wu et al. (2024); Chen et al. (2024). These benchmarks simulate the experimental design in psychological or cognitive research Wimmer and Perner (1983); Dyck et al. (2001), and contain elaborately crafted, story-based tasks that target ToM dimensions originally defined in psychology.
@@
Despite their wide coverage of ToM dimensions, these existing benchmarks all largely overlook the importance of comprehensive contextual understanding in human ToM. In daily lives, humans understand the minds of others based on long historical contexts of their personal backgrounds, rather than with only local circumstances on the spot. This is natural as human behaviors and mental states are also greatly influenced by global background factors of a person, such as social relationships, interactions with others, their personality and past experiences. However, the testing scenarios in existing ToM datasets are usually short of capturing these factors, as they are typically depicted with brief text pieces consisting of only superficial character actions and surrounding environment.
@@
In this paper, we propose CharToM-QA benchmark (“Charstands for “Character”) that evaluates the capability of LLMs on understanding ToM of characters in famous novel books. In CharToM-QA, the task takes the form of ToM-related question answering (QA) about characters within story plots. This setting naturally addresses the aforementioned challenges of most existing datasets due to the intrinsic features of story plots in novel books: 1) diverse social scenarios; 2) rich in complex social relationships and interactions; 3) high relevance to the whole book storyline. Thus, it alleviates heavy reliance on pre-determined rules to generate testing scenarios and raises higher requirements for comprehensively understanding context when evaluating ToM capability of current LLMs.
@@
For the dataset creation, we adopt an AI-assisted human annotation strategy that exploits publicly available book notes written by online reading app users. As users read, they can underline a specific text fragment and write their comments as book notes. These notes are often closely related to the selected fragments and their surrounding plots, containing valuable interpretation, analysis and thoughts of readers. We first collect a large number of user notes and their linked text from novel books. Then, as in most previous ToM assessments Nematzadeh et al. (2018); Sap et al. (2019); Tracey et al. (2022); Gandhi et al. (2023); Ma et al. (2023b); Wu et al. (2023), we keep the notes reflecting certain ToM dimensions of appointed characters, including belief, intention, emotion and desire. Based on these notes, GPT-4o is leveraged to generate ToM descriptions of the four dimensions about given characters. Afterwards, these descriptions are used to construct ToM-related questions with answers. Finally, the experiments are conducted in both generative and multichoice QA settings, where models are asked to make their responses given a question and corresponding story plot, as shown in Figure 1. During the whole process, strict validation procedures are carried out by experts with strong literature background.
@@
For evaluation, in generative QA, the qualities of responses are assessed based on the annotated answers. Particularly, traditional token-based metrics (e.g. Rouge Lin (2004)) and embedding-based metrics (e.g. cosine similarity between sentence embeddings) cannot give a fine-grained assessment that reflects how well a response matches the answer. Therefore, we design an evaluation protocol that mimics the process of grading papers. Specifically, we extract critical points from the answers, which are expected to be included in responses as bonus points. Then, GPT-4o is used as evaluator for assessment based on these bonus points. Besides, the evaluator also serves like a criticizer to point out defects existing in responses as penalty if there is any inappropriate statement. The bonus point coverage and penalty rate are used to comprehensively assess the qualities of model responses. In multichoice QA, we elaborately construct distraction choices and measure the model accuracy.
@@
We experiment with several current popular LLMs in both generative and multichoice QA settings, where different lengths of plot window are exposed to models. The experimental results show that GPT-4o consistently outperforms other LLMs across the four ToM dimensions in both settings. Moreover, human studies are conducted using multichoice QA. The results reveal that (1) Humans who have read the book greatly outperform the state-of-the-art LLM, while those who have not perform only on par with it. (2) Human ToM largely benefits from familiarity with novels. (3) Humans achieve higher accuracy after reading longer contexts. In contrast, the performances of LLMs almost stay stable with different lengths of plot window. This indicates that the ToM capability of current LLMs still struggles at dealing with complex scenarios and capturing nuanced historical contextual information for robust ToM comprehension.
@@ 2 Related Work
With the advent of LLMs, many benchmarks have been developed to evaluate the ToM capability of LLMs by simulating the cognitive experiments originally in psychology Grant et al. (2017); Nematzadeh et al. (2018); Sap et al. (2019); Le et al. (2019); Gandhi et al. (2023); Kosinski (2023); Wu et al. (2023); Kim et al. (2023); Gandhi et al. (2023); Zhou et al. (2023); Yu et al. (2024); Shapira et al. (2024); Xu et al. (2024a); Chan et al. (2024); Wu et al. (2024); Chen et al. (2024). Nematzadeh et al. (2018) proposed ToMi for evaluating the capability of models to reason about beliefs. Based on ToMi, Sap et al. (2019) introduced Social IQA that tests social and emotional intelligence of models. Wu et al. (2023) constructed Hi-ToM to assess higher-order ToM capability of models that involves recursive reasoning on others’ beliefs. Kim et al. (2023) build FANTOM to stress-test ToM within conversational contexts. Xu et al. (2024a) constructed OpenToM for assessing ToM with narrative stories containing explicit personality traits or preferences. Wu et al. (2024) created COKE consisting of cognitive chains that evaluate machine ToM capability to comprehend human activities. Chen et al. (2024) introduced ToMBench that encompasses 8 tasks and 31 abilities in social cognition.
Despite existing ToM benchmarks have covered most dimensions defined in psychology, the constructing procedures and testing scenarios of these datasets are still hindered by inherent limitations. 1) The story generation procedures of these benchmarks rely heavily on pre-determined rules and templates Nematzadeh et al. (2018); Le et al. (2019); Sap et al. (2019); Wu et al. (2023), constraining it from simulating diverse scenarios in reality. This also increases the risk of incorporating predictable regularities and spurious correlations into datasets, bringing about so-called “Clever Hans” phenomenon Lapuschkin et al. (2019). 2) The testing scenarios in these datasets are usually short of complex social relationships and interactions so that the importance of comprehensive contextual understanding is almost ignored. Appendix C shows a more detailed comparison of CharToM-QA benchmark with existing benchmarks.
3 Problem Formulation of ToM in Global Contexts
@@
In this paper, we consider a set of ToM dimensions prevalently studied in previous work in Section 2, Belief, Intention, Emotion and Desire. The importance of these dimensions in daily lives have been discussed in Apperly (2010). We follow the standard definitions of these dimensions in Apperly (2010), with the detailed definitions provided in Appendix B.
@@ 4 CharToM-QA: Assessing the Contextual Aspect of ToM
We create the CharToM-QA benchmark consisting of questions about the minds of appointed characters in novel stories. The benchmark is designed to assess the ToM capability of LLMs in understanding characters within global contexts. The definitions of the four studied ToM dimensions are given in Appendix B. The questions and answers are annotated by four human experts in literature, with an AI-assisted annotation strategy that leverages massive publicly available book notes from online reading platforms.
The dataset construction undergoes three steps: (1) book notes collection & filtering; (2) key note extraction & paraphrasing; (3) question generation & verification. Our dataset supports evaluation in both generative and multichoice QA settings.
@@ 4.1 Book Notes Collection & Filtering
Next, we filter the notes and keep those that can express any one of the four ToM dimensions of characters within story plots. The prompt used for note filtering is detailed in Appendix F. To ensure the validity of these filtered notes, we also manually check their correctness. Specifically, we sample 100 notes from four books with two annotators who are familiar with the books. For each note, the annotators are required to judge whether it appropriately describes the ToM of the appointed character. The results show that 1) 80% notes are judged as accurate; 2) 13% of the notes are open to subjective interpretation by different readers, yet they remain consistent with the book’s content and cannot be falsified; 3) only 7% notes are inappropriate, which demonstrates the overall validity of our filtered notes. Overall, 93% of notes can be treated as accurate.
@@ 4.2 Construction of Generative QA Task
At the extraction stage, the annotators are instructed to extract the critical part of a note that explicitly reflects one of the four ToM dimensions of a target character, which we call as “key note”. The basic principle of this stage is that annotators should only consider the literal meaning of a note with no additional inference. We elaborately craft examples for every ToM dimension and write a detailed manual to guide annotators. An annotator is only considered qualified after passing a labeling trial. They are given necessary information linked with a user note, including the target character, the underlined text and its surrounding context. After extraction, the key notes are kept for each ToM dimension. Nevertheless, the extracted key notes are often in a form of fragmented expressions that cannot be immediately used as complete ToM descriptions. Thus, a paraphrasing stage is required to convert fragmented parts into complete statements. For each note, we prompt GPT-4o to conduct paraphrasing based on key notes. The details of our used prompt template are presented in Appendix G. Annotators also check the paraphrased descriptions and make minor edit if necessary.
@@
In this step, we aim to construct ToM-related questions using those ToM descriptions. Figure 4 depicts the process of question generation. GPT-4o serves as a question generator to produce a set of candidate questions asking about target characters. During generation, along with the ToM description, we also provide GPT-4o with corresponding underlined text and its surrounding context of current story plot. For each ToM description, GPT-4o is asked to generate several candidate questions that adhere to the following requirements:
@@
Subsequently, a verification procedure is involved to preliminarily filter out inappropriate candidates that violate the above requirements. We find that it is better not to let GPT-4o be aware of the story plot. This is because GPT-4o tends to build trivial connections between question and answer using any possible detail in the story regardless of importance, preserving many low-quality candidates. The prompt templates used for generation and verification are given in Appendix H. In this way, we keep at most 4 candidate questions for each ToM description. Then, annotators are responsible for choosing the best one out of these candidates, where slight modifications by human are allowed during examination.
@@ 4.3 Construction of Multichoice QA Task
Besides the aformentioned Generative QA setting, we support multichoice QA evaluation with elaborately constructed distraction choices for each question. Given the question-answer pair within a story plot, GPT-4o is used to produce distraction choices that are seemingly plausible but actually unreasonable. Appendix J provides the details of constructing such difficult distraction choices. Finally, a question is paired with four choices including the correct answer and three incorrect ones, as shown in the Appendix M.
@@ 5 Evaluation Protocols for Generative QA
Until now, we have acquired ToM-related question-answer pairs along with their linked story plots, which can be used to conduct evaluation in a generative QA setting. Models are given story plots and asked questions to make responses whose qualities are assessed using corresponding answers as reference. Conventionally, token-based metrics (e.g. Rouge Lin (2004)) and embedding-based metrics (e.g. cosine similarity between sentence embeddings) are mainly used for assessing generative QA tasks. However, these two types of metrics have their inherent limitations. For token-based metrics, they cannot deal with the cases where 1) same meanings are expressed with different tokens; 2) non-critical tokens dominate in a sentence. Meanwhile, embedding-based metrics only coarsely measure the semantic similarity but cannot explicitly indicate the pros and cons of a response.
Therefore, inspired by the process of grading papers, we design an evaluation protocol that inspects the bonus points and penalty of a model response. Specifically, for each question, we extract the critical points of its reference answer with the assistance of GPT-4o, which are expected to be included as bonus points in a response. The core requirements for bonus point extraction are: 1) Bonus points must be derived from reference answers with no hallucination; 2) Different bonus points should orient to different aspects of an answer. The statistics about the bonus points of questions in each dimension is given in Appendix D. During evaluation, a GPT-4o evaluator measures the coverage of bonus points as an indicator of response quality. Moreover, it also criticizes the defects in a response as penalty if there is any inappropriate statement. In this way, both bonus point coverage and penalty rate are used as a comprehensive assessment of response quality. The results in Appendix E also demonstrate that our designed evaluation protocol is more correlated with human judgments than conventional metrics. Particularly, the prompts for extracting bonus points and conducting such evaluation protocol are detailed in Appendices I and L, respectively. Appendix M gives illustrative cases of such evaluation protocol.